40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 5 11:25:21 2015, ericq, Update, SUS, ETMY OL laser dead ETMY-OLSUM.png
    Reply  Mon Oct 5 16:41:22 2015, Steve, Update, SUS, ETMY OL laser replaced 
Message ID: 11667     Entry time: Mon Oct 5 11:25:21 2015     Reply to this: 11668
Author: ericq 
Type: Update 
Category: SUS 
Subject: ETMY OL laser dead 

Gautam alerted me that the Y arm looked like it was being dithered, even though the ASS was turned off. I found that the ETMY OL signals were garbage, leading to the servos flipping back and forth between their rails. 

We went out to the ETMY table, and found the HeNe laser to be emitting a paltry <0.5mW; the OL QPD could not register the puny beam incident on it.

Here is the last 30 days of OL_SUM:

Steve will replace the laser this afternoon. 

ELOG V3.1.3-